OpenStudy (anonymous):

Two ribbons are selected at random from a container holding 5 purple and 6 white ribbons. Find the probability that both ribbons are white.

OpenStudy (agent0smith):

First you need the total number of ribbons. The probability of getting two ribbons (this time will be without replacement) will be P(1st ribbon)*P(2nd ribbon) P(1st ribbon) is just the number of white ribbons, over the total. P(2nd ribbon) is the number of remaining white ribbons, over the remaining total ribbons.

OpenStudy (anonymous):

6/11 & 5/11 ? or is this assuming that I'm removing both white ribbons? (Sorry, I make math more complicated for myself by over thinking everything.)

OpenStudy (anonymous):

Wait, >.< 6/11 then ... 5/10 = 1/2 ? (See what I mean. )

OpenStudy (agent0smith):

haha, excellent. You found your mistake and corrected yourself! :)

OpenStudy (agent0smith):

Now you just need: P(1st ribbon)*P(2nd ribbon)

OpenStudy (anonymous):

Do I just multiply them? Or, am I thinking incorrectly again? lol.

OpenStudy (agent0smith):

Don't doubt yourself, you're correct. You can also use combinations here, if you've done those: from 6 white ribbons, choose 2, over from 11 total ribbons choose 2. \[\Large \frac{ 6C2 }{ 11C2 }\]

OpenStudy (anonymous):

I'm not sure what to do with that equation, I haven't worked with them.

OpenStudy (agent0smith):

Never mind it then :)

OpenStudy (agent0smith):

Just use P(1st ribbon)*P(2nd ribbon)

OpenStudy (anonymous):

6/11*1/2

OpenStudy (anonymous):

or 3/11 :)

OpenStudy (agent0smith):

Yes :)